OpenStudy (anonymous):

write an equation in the form y=mx+b for the line which passes through the points (0,2) and(4,3)

OpenStudy (godlovesme):

do u know how to calculate the slope using the given ordered pairs?

OpenStudy (godlovesme):

@ccf122 u there?

OpenStudy (anonymous):

im back sorry

OpenStudy (godlovesme):

it's okay :) so do u know how to calculate the slope using the given ordered pairs?

OpenStudy (anonymous):

yes give me a min

OpenStudy (godlovesme):

okiee C:

OpenStudy (anonymous):

i dont know bout this answer but it gave me 0.25

OpenStudy (godlovesme):

did u use online calculator?

OpenStudy (anonymous):

yep

OpenStudy (godlovesme):

can i show u how to do it without one? :)

OpenStudy (anonymous):

sure

OpenStudy (godlovesme):

okiee :)) so we have two ordered pairs (0,2) and (4,3) ordered pairs are represented as (x,y) so in this case our Xs are 0 and 4 out Ys are 2 and 3 confusing?

OpenStudy (anonymous):

no and ik to subtact both and divide the fraction that come with it

OpenStudy (anonymous):

i see what i did

OpenStudy (godlovesme):

what do u mean by "subtract both and divide the fraction that come with it" ?

OpenStudy (anonymous):

how i did it was 3-2=1 and 4-0=4 which is 1/4 divid and the answer i goy was 0.25

OpenStudy (godlovesme):

woow smart :D so are u familiar with the y=mx+b formula?

OpenStudy (anonymous):

no not realy

OpenStudy (godlovesme):

okay so y=mx+b is known as the slope-intercept formula m represents the slope b represents the y-intercept

OpenStudy (anonymous):

ok so which y are we going to do

OpenStudy (godlovesme):

y-intercept is where the line hits the y-axis If we put it in an ordered pair form the x MUST be 0 so in this case our y-intercept is?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

4 im sorry very tired to let you know

OpenStudy (godlovesme):

not quite, the ordered pair should look something like this (0, y) and y can be any number but x MUST be 0

OpenStudy (anonymous):

so are we just using 0 and 2

OpenStudy (godlovesme):

yes which means the y-intercept is?

OpenStudy (godlovesme):

is X is 0 y is? :P

OpenStudy (godlovesme):

when X **

OpenStudy (godlovesme):

y= 1/4 x + 2 :P

OpenStudy (godlovesme):

@ccf122 hope that helped! :)

OpenStudy (anonymous):

thank god for u